# Reasoning Beyond Limits: Advances and Open Problems for LLMs

### v1

@@ IV-W Reflective Augmentation (RefAug)
Zhang et al. [126] propose Reflective Augmentation (RefAug), a novel technique designed to embed problem reflection into training data for language models, enhancing their problem-solving abilities on mathematical and coding tasks. Unlike conventional data augmentation that merely increases training volume, RefAug integrates alternative reasoning approaches, analogies, and diverse perspectives alongside each problem and its solution, thereby training models to engage in both forward and reflective reasoning. Extensive experiments demonstrate that this method improves accuracy in math tasks by 6.8 points, boosts code performance by 3.5 points in Pass@1, and achieves an error correction rate of 81. 11%, outperforming traditional chain-of-threat methods.
TABLE VI: Challenges in Enhancing LLM Reasoning
@@
Topic Core Issue Proposed Approach Observed Outcomes Overall Significance Improving Reasoning Without Human Supervision Enhancing multi-step reasoning without human annotations. OmegaPRM [26] – a divide-and-conquer Monte Carlo Tree Search that autonomously collects process supervision data. Boosts Gemini Pro’s success on MATH500 from 51% to 69.4% and on GSM8K from 86.4% to 93.6%. Demonstrates a scalable, cost-effective method for self-supervised reasoning improvement. Synthetic Data for Math Olympiad LLMs show a reasoning gap in chained, multi-hop math problems. ToRA [127] and MuMath-Code [128] combine synthetic data generation with tool integration. ToRA-7B scores 44.6% on MATH; MuMath-Code-70B achieves 90.7% on GSM8K and 55.1% on MATH. Highlights the potential of data augmentation and external tool use to bridge reasoning gaps. Criticality of Online Data for RLHF Offline datasets often lack diversity for optimal RLHF convergence. Comparing online RL methods (e.g., PPO) and introducing Hybrid Preference Optimization (HyPO) [129]. Online approaches require only partial coverage and outperform offline methods. Emphasizes that on-policy data is crucial for effective model alignment. Multilingual RLHF Transfer RLHF techniques predominantly focus on English, limiting performance in other languages. Generating synthetic multilingual feedback and mixing language data during training [130]. Training on English-only data yields up to 7% improvements in other languages; multi-language data boosts win rates by up to 19%. Demonstrates significant cross-lingual transfer, enhancing global model alignment. DPO for LLM as a Judge Standard reward models struggle with robust auto-evaluation. SFR-Judges [131] integrate Direct Preference Optimization (DPO) with supervised fine-tuning loss using chain-of-thought critiques. Achieves  44% performance on Alpaca Eval and  81% pairwise comparison accuracy. Improves generative judge capabilities and downstream model evaluation. SFT Memorization vs. RL Generalization SFT tends to overfit, hindering out-of-domain generalization. Reinforcement Fine-Tuning (RFT) [132] applied on top of an SFT baseline. Yields up to 61% improvement on out-of-domain tasks and +34% on visual navigation tasks. Underlines the need to combine SFT with RL for robust model adaptation. Impact of Structured Prompting Rigid formats (e.g., JSON, XML) may restrict reasoning ability. Comparing constraint decoding, format-restricting instructions, and two-step approaches [133]. Mixed outcomes: some models perform optimally with JSON; others see performance drops. Highlights the trade-off between output consistency and cognitive flexibility. Effectiveness of Long-Context RAG LLMs degrade when processing contexts beyond optimal lengths. Systematic evaluation of retrieval thresholds and failure modes [134, 135]. Optimal performance at  32k–64k tokens, with degradation and issues (e.g., repetition) beyond. Suggests a need for improved long-context training and retrieval strategies. Role of the Reference Model in DPO The strength of the reference model affects DPO alignment quality. Experimenting with varied KL constraints and using stronger, compatible reference models [136]. Performance increases from 12.36 to 20.25 with optimal settings. Underlines that a well-chosen reference model is vital for DPO success. Online vs. Offline Alignment Offline methods excel in classification but lag in generative tasks. Empirical comparison of online RLHF (e.g., PPO, RLOO) versus offline contrastive approaches [71]. Online methods achieve higher win rates, despite offline models performing better in pairwise tasks. Highlights the value of on-policy sampling for generative quality and the trade-offs in alignment. Leveraging LLMs for Code Compilation LLMs have untapped potential for code compilation and optimization tasks. LLM Compiler [137]: Fine-tuning Code Llama on large corpora of LLVM IR and assembly code. Achieves 77% of autotuning potential, 45% disassembly accuracy, and 14% exact match. Offers a scalable, cost-effective foundation for compiler research and optimization. Training MoE Architectures Efficiently training large-scale MoE models poses communication challenges. Leveraging advanced communication libraries (e.g., DeepSpeed, MegatronLM) for MoE [138, 139, 140, 141]. Enables development of DeepSeek-R1 with efficient training (approx. \$6M cost) and optimized inference. Advances efficient MoE parallelism, crucial for scaling large language models.
V Challenges and Open problems
@@
This section presents a discussion of key challenges in advancing LLM capabilities. It highlights the potential of autonomous methods to enhance multi-step reasoning without human intervention while also noting the persistent difficulties in solving chained, multi-hop problems. It discusses the benefits and limitations of reinforcement learning from human feedback, particularly in multilingual settings, and examines the trade-offs between structured prompting and maintaining reasoning flexibility. In addition, it addresses the challenges associated with long-context retrieval and the opportunities in leveraging LLMs for code compilation and efficient Mixture-of-Experts architectures. Table VI presents the challenges in enhancing LLM reasoning.
V-A Can we improve Reasoning without human supervision?
@@
Recent advances by Google DeepMind [26] demonstrate that autonomous approaches can effectively enhance multi-step reasoning in large language models. This work introduces OmegaPRM, a novel divide-and-conquer Monte Carlo Tree Search algorithm that autonomously collects high-quality process supervision data to train Process Reward Models (PRMs) and significantly boost reasoning performance. OmegaPRM employs a binary search strategy to identify the first error in a chain-of-thought, isolating mistakes and balancing positive and negative examples, which enables the generation of over 1.5 million supervision annotations without any human intervention. These annotations are used to train a PRM with a pointwise soft label objective, and when combined with weighted self-consistency, the method dramatically improves performance—boosting the instruction-tuned Gemini Pro model’s success rate on MATH500 from 51% to 69.4% and on GSM8K from 86.4% to 93.6%, while similarly enhancing Gemma2 27B’s outcomes. Overall, OmegaPRM offers a cost-effective, scalable solution that reduces computational overhead and refines LLM reasoning by rewarding intermediate progress, all achieved without direct human supervision.
V-B Will Synthetic Data Win the AI Math Olympiad?
@@
The study by Hosseini et al. [142] delves into the grade-school math problem-solving abilities of LLMs by examining their performance on paired problems—where the solution to the first is essential for resolving the second—to uncover what the authors term a ”reasoning gap.” While many LLMs perform adequately on isolated math questions, their accuracy drops significantly when tasked with chained, multi-hop problems, revealing that the contextual overload and additional reasoning steps introduce challenges not apparent in standard benchmarks. Notably, this gap is more evident in smaller, cost-efficient, and math-specialized models, suggesting that even targeted fine-tuning can lead to overfitting and hinder generalization to composite tasks. The paper also highlights that improvements from instruction-tuning and code generation vary with model size, and that the reasoning shortfall is attributable not to data leakage but to inherent distractions and insufficient second-hop reasoning capabilities. Overall, the findings prompt a reevaluation of how LLMs are assessed for complex, multi-step problem solving, urging the development of more nuanced benchmarks that capture true reasoning proficiency.
@@
Gou et al. [127] presents ToRA, a series of Tool-integrated Reasoning Agents that enhance large language models’ ability to tackle complex mathematical problems by integrating natural language reasoning with external computational tools such as symbolic solvers and computation libraries. Recognizing that traditional LLMs struggle with intricate mathematical tasks due to limitations in pure language processing, the authors propose a novel training regimen that involves curating interactive tool-use trajectories from mathematical datasets, leveraging imitation learning on annotated data, and applying output space shaping to fine-tune reasoning behavior. Empirical evaluations demonstrate that ToRA models outperform existing open-source alternatives on ten mathematical reasoning benchmarks, yielding average absolute improvements between 13% and 19%. Notably, the ToRA-7B model achieves a score of 44.6% on the challenging MATH dataset, surpassing the previous best open-source model by 22 percentage points, while ToRA-Code-34B becomes the first open-source model to exceed 50% accuracy on MATH—outperforming GPT-4’s chain-of-thought approach and rivaling its program-based problem solving.
@@
Yin et al. [128] proposes a novel approach that bridges two previously distinct research directions in mathematical reasoning for LLMs: external tool integration and data augmentation. Traditionally, tool-use LLMs enhance performance by interfacing with Python interpreters, while tool-free methods rely on enriching math datasets. To harness the benefits of both, the authors introduce MuMath-Code, which fine-tunes Llama-2 on an augmented dataset generated via multi-perspective data augmentation and synthesized code-nested solutions. The model employs a two-stage training strategy: Stage-1 involves fine-tuning on chain-of-thought data to establish robust reasoning patterns, and Stage-2 further refines the model on the newly augmented code-nested data. At inference, MuMath-Code generates executable code that interacts with an external Python interpreter to produce precise computation results. The approach is validated through extensive experiments, with the MuMath-Code-7B model achieving scores of 83.8 on GSM8K and 52.4 on MATH, and the MuMath-Code-70B model setting new state-of-the-art performance among open-source methods—90.7% on GSM8K and 55.1% on MATH. These results underscore the effectiveness of combining tool use with data augmentation and highlight a promising direction for future research in enhancing mathematical reasoning in LLMs.
V-C How critical is online data for RLHF?
@@
Song et al. [129] presents a rigorous theoretical and empirical comparison between online reinforcement learning (RL) methods and offline contrastive techniques for fine-tuning large language models using human preference data. By introducing the notion of dataset coverage—a measure of how well the training data represents the test distribution—the authors establish that while offline methods like Direct Preference Optimization (DPO) require a stringent global coverage condition for optimal convergence, online RL methods, such as Proximal Policy Optimization (PPO), only need a weaker partial coverage condition. This theoretical distinction helps explain why online methods often outperform their offline counterparts, particularly in scenarios where the offline dataset lacks sufficient diversity. Motivated by these insights, the paper introduces Hybrid Preference Optimization (HyPO), a novel algorithm that leverages offline contrastive optimization alongside online data for KL regularization. The hybrid approach achieves superior performance compared to pure offline methods and retains computational and memory efficiency. I find PPO most compelling among the online RLHF methods discussed—PPO, RLOO, Online-DPO, ReMax, and J-BOND—given its robust theoretical foundation and proven success in dynamically generating and integrating new data during the training process.
V-D Does RLHF transfer to different languages?
@@
Dang et al. [130] presents a comprehensive study on extending preference optimization techniques for LLMs to a multilingual setting, addressing the gap in current research that predominantly focuses on English and Chinese. The authors propose a novel and scalable method to generate high-quality multilingual feedback data, which balances data coverage and facilitates cross-lingual transfer during training. By creating synthetic multilingual preference datasets—using approximately 50K English prompts translated into 22 languages—and designing various dataset mixtures, the study rigorously evaluates both offline (DPO) and online (RLOO) RLHF methods. Key findings reveal that even training on English-only data can yield up to 7% improvements in other languages, while incorporating multiple languages can boost win rates by as much as 19% on unseen languages. Notably, online RLHF (RLOO) consistently outperforms its offline counterpart by up to 10.6%, demonstrating superior transfer capabilities. The resulting preference-trained model achieves a 54.4% win-rate against the current state-of-the-art Aya 23 8B and outperforms several widely used models, thereby significantly advancing multilingual LLM alignment across 23 languages and impacting nearly half of the world’s population.
V-E Can DPO improve LLM as a Judge?

### v2

@@ VI-W Reflective Augmentation (RefAug)
Zhang et al. [zhang2024learn] propose Reflective Augmentation (RefAug), a novel technique designed to embed problem reflection into training data for language models, enhancing their problem-solving abilities on mathematical and coding tasks. Unlike conventional data augmentation that merely increases training volume, RefAug integrates alternative reasoning approaches, analogies, and diverse perspectives alongside each problem and its solution, thereby training models to engage in both forward and reflective reasoning. Extensive experiments demonstrate that this method improves accuracy in math tasks by 6.8 points, boosts code performance by 3.5 points in Pass@1, and achieves an error correction rate of 81. 11%, outperforming traditional chain-of-threat methods.
TABLE VIII: Challenges in Enhancing LLM Reasoning
@@
Topic Core Issue Proposed Approach Observed Outcomes Overall Significance Improving Reasoning Without Human Supervision Enhancing multi-step reasoning without human annotations. OmegaPRM [luo2024improve] – a divide-and-conquer Monte Carlo Tree Search that autonomously collects process supervision data. Boosts Gemini Pro’s success on MATH500 from 51% to 69.4% and on GSM8K from 86.4% to 93.6%. Demonstrates a scalable, cost-effective method for self-supervised reasoning improvement. Synthetic Data for Math Olympiad LLMs show a reasoning gap in chained, multi-hop math problems. ToRA [gou2023tora] and MuMath-Code [yin2024mumath] combine synthetic data generation with tool integration. ToRA-7B scores 44.6% on MATH; MuMath-Code-70B achieves 90.7% on GSM8K and 55.1% on MATH. Highlights the potential of data augmentation and external tool use to bridge reasoning gaps. Criticality of Online Data for RLHF Offline datasets often lack diversity for optimal RLHF convergence. Comparing online RL methods (e.g., PPO) and introducing Hybrid Preference Optimization (HyPO) [song2024importance]. Online approaches require only partial coverage and outperform offline methods. Emphasizes that on-policy data is crucial for effective model alignment. Multilingual RLHF Transfer RLHF techniques predominantly focus on English, limiting performance in other languages. Generating synthetic multilingual feedback and mixing language data during training [dang2024rlhf]. Training on English-only data yields up to 7% improvements in other languages; multi-language data boosts win rates by up to 19%. Demonstrates significant cross-lingual transfer, enhancing global model alignment. DPO for LLM as a Judge Standard reward models struggle with robust auto-evaluation. SFR-Judges [wang2024direct] integrate Direct Preference Optimization (DPO) with supervised fine-tuning loss using chain-of-thought critiques. Achieves  44% performance on Alpaca Eval and  81% pairwise comparison accuracy. Improves generative judge capabilities and downstream model evaluation. SFT Memorization vs. RL Generalization SFT tends to overfit, hindering out-of-domain generalization. Reinforcement Fine-Tuning (RFT) [chu2025sftmemorizesrlgeneralizes] applied on top of an SFT baseline. Yields up to 61% improvement on out-of-domain tasks and +34% on visual navigation tasks. Underlines the need to combine SFT with RL for robust model adaptation. Impact of Structured Prompting Rigid formats (e.g., JSON, XML) may restrict reasoning ability. Comparing constraint decoding, format-restricting instructions, and two-step approaches [tam2024let]. Mixed outcomes: some models perform optimally with JSON; others see performance drops. Highlights the trade-off between output consistency and cognitive flexibility. Effectiveness of Long-Context RAG LLMs degrade when processing contexts beyond optimal lengths. Systematic evaluation of retrieval thresholds and failure modes [leng2024long, laban2024summary]. Optimal performance at  32k–64k tokens, with degradation and issues (e.g., repetition) beyond. Suggests a need for improved long-context training and retrieval strategies. Role of the Reference Model in DPO The strength of the reference model affects DPO alignment quality. Experimenting with varied KL constraints and using stronger, compatible reference models [liu2024understanding]. Performance increases from 12.36 to 20.25 with optimal settings. Underlines that a well-chosen reference model is vital for DPO success. Online vs. Offline Alignment Offline methods excel in classification but lag in generative tasks. Empirical comparison of online RLHF (e.g., PPO, RLOO) versus offline contrastive approaches [tang2024understanding]. Online methods achieve higher win rates, despite offline models performing better in pairwise tasks. Highlights the value of on-policy sampling for generative quality and the trade-offs in alignment. Leveraging LLMs for Code Compilation LLMs have untapped potential for code compilation and optimization tasks. LLM Compiler [cummins2024meta]: Fine-tuning Code Llama on large corpora of LLVM IR and assembly code. Achieves 77% of autotuning potential, 45% disassembly accuracy, and 14% exact match. Offers a scalable, cost-effective foundation for compiler research and optimization. Training MoE Architectures Efficiently training large-scale MoE models poses communication challenges. Leveraging advanced communication libraries (e.g., DeepSpeed, MegatronLM) for MoE [deepspeed-moe, megatron-lm, llm-foundry, deepep]. Enables development of DeepSeek-R1 with efficient training (approx. \$6M cost) and optimized inference. Advances efficient MoE parallelism, crucial for scaling large language models.
VII Challenges and Open problems
@@
This section presents a discussion of key challenges in advancing LLM capabilities. It highlights the potential of autonomous methods to enhance multi-step reasoning without human intervention while also noting the persistent difficulties in solving chained, multi-hop problems. It discusses the benefits and limitations of reinforcement learning from human feedback, particularly in multilingual settings, and examines the trade-offs between structured prompting and maintaining reasoning flexibility. In addition, it addresses the challenges associated with long-context retrieval and the opportunities in leveraging LLMs for code compilation and efficient Mixture-of-Experts architectures. Table VIII presents the challenges in enhancing LLM reasoning.
VII-A Can we improve Reasoning without human supervision?
@@
Recent advances by Google DeepMind [luo2024improve] demonstrate that autonomous approaches can effectively enhance multi-step reasoning in large language models. This work introduces OmegaPRM, a novel divide-and-conquer Monte Carlo Tree Search algorithm that autonomously collects high-quality process supervision data to train Process Reward Models (PRMs) and significantly boost reasoning performance. OmegaPRM employs a binary search strategy to identify the first error in a chain-of-thought, isolating mistakes and balancing positive and negative examples, which enables the generation of over 1.5 million supervision annotations without any human intervention. These annotations are used to train a PRM with a pointwise soft label objective, and when combined with weighted self-consistency, the method dramatically improves performance boosting the instruction-tuned Gemini Pro model’s success rate on MATH500 from 51% to 69.4% and on GSM8K from 86.4% to 93.6%, while similarly enhancing Gemma2 27B’s outcomes. Overall, OmegaPRM offers a cost-effective, scalable solution that reduces computational overhead and refines LLM reasoning by rewarding intermediate progress, all achieved without direct human supervision.
This emerging paradigm opens new research directions for optimizing automated supervision signals. One promising avenue is integrating uncertainty estimation or probabilistic error localization into the reward signal design. Furthermore, combining self-supervised distillation with programmatic supervision, such as that derived from automated theorem provers or formal verifiers, could reduce reliance on manually labeled reasoning trajectories. Investigating how these methods generalize across domains and reasoning tasks (e.g., symbolic, scientific, legal) remains a key challenge for future work.
VII-B Will Synthetic Data Win the AI Math Olympiad?
@@
The study by Hosseini et al. [hosseini2024not] delves into the grade-school math problem-solving abilities of LLMs by examining their performance on paired problems where the solution to the first is essential for resolving the second to uncover what the authors term a ”reasoning gap.” While many LLMs perform adequately on isolated math questions, their accuracy drops significantly when tasked with chained, multi-hop problems, revealing that the contextual overload and additional reasoning steps introduce challenges not apparent in standard benchmarks. Notably, this gap is more evident in smaller, cost-efficient, and math-specialized models, suggesting that even targeted fine-tuning can lead to overfitting and hinder generalization to composite tasks. The paper also highlights that improvements from instruction-tuning and code generation vary with model size, and that the reasoning shortfall is attributable not to data leakage but to inherent distractions and insufficient second-hop reasoning capabilities. Overall, the findings prompt a reevaluation of how LLMs are assessed for complex, multi-step problem solving, urging the development of more nuanced benchmarks that capture true reasoning proficiency.
@@
Gou et al. [gou2023tora] presents ToRA, a series of Tool-integrated Reasoning Agents that enhance large language models’ ability to tackle complex mathematical problems by integrating natural language reasoning with external computational tools such as symbolic solvers and computation libraries. Recognizing that traditional LLMs struggle with intricate mathematical tasks due to limitations in pure language processing, the authors propose a novel training regimen that involves curating interactive tool-use trajectories from mathematical datasets, leveraging imitation learning on annotated data, and applying output space shaping to fine-tune reasoning behavior. Empirical evaluations demonstrate that ToRA models outperform existing open-source alternatives on ten mathematical reasoning benchmarks, yielding average absolute improvements between 13% and 19%. Notably, the ToRA-7B model achieves a score of 44.6% on the challenging MATH dataset, surpassing the previous best open-source model by 22 percentage points, while ToRA-Code-34B becomes the first open-source model to exceed 50% accuracy on MATH outperforming GPT-4’s chain-of-thought approach and rivaling its program-based problem solving.
@@
Yin et al. [yin2024mumath] proposes a novel approach that bridges two previously distinct research directions in mathematical reasoning for LLMs: external tool integration and data augmentation. Traditionally, tool-use LLMs enhance performance by interfacing with Python interpreters, while tool-free methods rely on enriching math datasets. To harness the benefits of both, the authors introduce MuMath-Code, which fine-tunes Llama-2 on an augmented dataset generated via multi-perspective data augmentation and synthesized code-nested solutions. The model employs a two-stage training strategy: Stage-1 involves fine-tuning on chain-of-thought data to establish robust reasoning patterns, and Stage-2 further refines the model on the newly augmented code-nested data. At inference, MuMath-Code generates executable code that interacts with an external Python interpreter to produce precise computation results. The approach is validated through extensive experiments, with the MuMath-Code-7B model achieving scores of 83.8 on GSM8K and 52.4 on MATH, and the MuMath-Code-70B model setting new state-of-the-art performance among open-source methods 90.7% on GSM8K and 55.1% on MATH. These results underscore the effectiveness of combining tool use with data augmentation and highlight a promising direction for future research in enhancing mathematical reasoning in LLMs.
These findings motivate several open research directions. One important avenue involves building curriculum-based synthetic datasets that gradually increase reasoning complexity to improve generalization across math tasks. Another promising line of work is the design of adaptive tool-augmented agents capable of deciding when and how to invoke external solvers during problem solving. Further research could also investigate the cross-domain transferability of these synthetic math reasoning skills to adjacent fields such as symbolic logic, formal verification, or algorithmic coding tasks.
VII-C How critical is online data for RLHF?
@@
Song et al. [song2024importance] presents a rigorous theoretical and empirical comparison between online reinforcement learning (RL) methods and offline contrastive techniques for fine-tuning large language models using human preference data. By introducing the notion of dataset coverage a measure of how well the training data represents the test distribution the authors establish that while offline methods like Direct Preference Optimization (DPO) require a stringent global coverage condition for optimal convergence, online RL methods, such as Proximal Policy Optimization (PPO), only need a weaker partial coverage condition. This theoretical distinction helps explain why online methods often outperform their offline counterparts, particularly in scenarios where the offline dataset lacks sufficient diversity. Motivated by these insights, the paper introduces Hybrid Preference Optimization (HyPO), a novel algorithm that leverages offline contrastive optimization alongside online data for KL regularization. The hybrid approach performs better than pure offline methods and retains computational and memory efficiency. I find PPO most compelling among the online RLHF methods discussed PPO, RLOO, Online-DPO, ReMax, and J-BOND given its robust theoretical foundation and proven success in dynamically generating and integrating new data during the training process.
This line of work opens several promising research directions. One direction involves the design of adaptive data collection strategies that prioritize coverage improvement during online training. Additionally, future work could explore principled hybrid architectures that dynamically switch between offline and online regimes based on distributional drift or model uncertainty. Another area of interest is developing new preference modeling paradigms that extend beyond binary or scalar reward signals to support structured or multi-dimensional human feedback. Finally, benchmarking these approaches across diverse instruction-following and dialog tasks can help determine their real-world robustness and generalization.
VII-D Does RLHF transfer to different languages?
@@
Dang et al. [dang2024rlhf] present a comprehensive study on extending preference optimization techniques for LLMs to a multilingual setting, addressing the gap in current research that predominantly focuses on English and Chinese. The authors propose a novel and scalable method to generate high-quality multilingual feedback data, which balances data coverage and facilitates cross-lingual transfer during training. By creating synthetic multilingual preference datasets using approximately 50K English prompts translated into 22 languages and designing various dataset mixtures, the study rigorously evaluates both offline (DPO) and online (RLOO) RLHF methods. Key findings reveal that even training on English-only data can yield up to 7% improvements in other languages, while incorporating multiple languages can boost win rates by as much as 19% on unseen languages. Notably, online RLHF (RLOO) consistently outperforms its offline counterpart by up to 10.6%, demonstrating superior transfer capabilities. The resulting preference-trained model achieves a 54.4% win-rate against the current state-of-the-art Aya 23 8B and outperforms several widely used models, thereby significantly advancing multilingual LLM alignment across 23 languages and impacting nearly half of the world’s population.
These results highlight a promising frontier for cross-lingual alignment via preference learning. Future research could explore more granular multilingual signal integration, such as culturally adaptive reward models or region-specific annotation schemes. Another direction involves dynamically adjusting language mixing ratios during training based on feedback quality or model confidence. In addition, aligning multilingual LLMs using code-switching inputs or zero-shot preference extrapolation may offer insights into generalized policy learning across languages and dialects. Finally, evaluating multilingual alignment performance across low-resource and typologically diverse languages remains a crucial step toward equitable LLM development.
VII-E Can DPO improve LLM as a Judge?